# Fundamental multiplication formulas

### v1

@@ 4.3. The geometric definition of the standard basis
(2) If $A\in\overset{\circ\!\circ\!\bullet}{\Xi}$ then, by Corollary 2.7, $W_{\lambda}={\check{W}}_{\lambda^{+}}$ and ${\check{W}}_{\mu}={\check{W}}_{\mu^{\bullet}}$.
@@
If $d_{A}s_{r}=s_{r}d_{A}$ then, by Lemma 2.2 (2), $a_{n+1,n+1}>0$. Thus, $A\in\overset{\bullet}{\Xi}$ and one of the two disjoint sets in (4.12.1) is a double cosets in ${\check{W}}$, depending on $d_{A}\in{\check{W}}$ or not.
@@
Recalled the standard basis $\{\phi_{\alpha,\beta}^{d}\mid(\alpha,d,\beta)\in\check{\mathcal{D}}(n,r)\}$ for $S_{{\boldsymbol{q}}}^{\textsc{d}}(n,r)$ given in Corollary 4.3. With the bijection $\eta$ in Lemma 4.12, we may also label the basis element by signed matrix notation:
@@
(1) For $A\in\overset{\circ\!\circ\!\circ}{\Xi}$ and ${\epsilon}\in\{+,-\}$, if $A$ defines ${}^{+}{A}^{{\epsilon}},^{-}{A}^{-{\epsilon}}\in\overset{\circ\!\circ\!\circ}{\Xi}_{+}\sqcup\overset{\circ\!\circ\!\circ}{\Xi}_{-}$, then $\eta(^{+}{A}^{{\epsilon}})=(\lambda^{+},d_{1},\mu^{\epsilon})$ and $\eta(^{-}{A}^{-{\epsilon}})=(\lambda^{-},d_{2},\mu^{-{\epsilon}})$ (see (6.1.1) for notational convention and Definition 4.11 for the definition of $d_{1},d_{2}$) which define distinct standard basis elements $\phi_{{}^{+}{A}^{{\epsilon}}}$ and $\phi_{{}^{-}{A}^{-{\epsilon}}}$ according their row or column weights given in (4.12.5). This distinction can be reflected by the distinction of the ${\mathsf{m}}_{1}\times{\mathsf{m}}_{2}$ subset matrices $\big{(}{\check{R}}_{i}^{\lambda^{+}}\cap d_{1}{\check{R}}_{j}^{\mu^{\epsilon}}\big{)}$ and $\big{(}{\check{R}}_{i}^{\lambda^{-}}\cap d_{2}{\check{R}}_{j}^{\mu^{-{\epsilon}}}\big{)}$ since
@@
(2) For $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$, we also have ${}^{+}{A}^{{\epsilon}},^{-}{A}^{-{\epsilon}}\in\overset{\bullet\!\circ\!\bullet}{\Xi}_{+}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}_{-}$. In this case, the standard basis elements $\phi_{{}^{+}{A}^{{\epsilon}}}$ and $\phi_{{}^{-}{A}^{-{\epsilon}}}$ have the same domain and codomain, but are defined by distinct double cosets ${\check{W}}_{\lambda^{\bullet}}d_{1}{\check{W}}_{\mu^{\bullet}}$ and ${\check{W}}_{\lambda^{\bullet}}d_{2}{\check{W}}_{\mu^{\bullet}}$. Of course, the conjugate intersections ${\check{W}}_{\lambda^{\bullet}}\cap({\check{W}}_{\mu^{\bullet}})^{d_{1}^{-1}}$ and ${\check{W}}_{\lambda^{\bullet}}\cap({\check{W}}_{\mu^{\bullet}})^{d_{2}^{-1}}$ are also distinct in this case.
@@
The theorem shows that the ${\mathcal{A}}$-algebra $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ is the “quantumization” (in the sense of [8, p.17]) of the algebras $\{\mathfrak{F}_{{\mathrm{SO}}_{2r}(q)}({\mathcal{X}}\times{\mathcal{X}})\}_{q\in\mathcal{P}}$ which is the same as the algebra defined in [14, (5)]. It should be point out that the dimension formula #$\#\Xi_{\textsc{d}}$ given in [14, Lem. 4.2.1] is incorrect as the definition of the index set $\Xi_{\textsc{d}}=\Xi^{+}\sqcup\Xi^{0}\sqcup\Xi^{-}$ mistakenly used $\Xi^{0}$ instead of $\overset{\bullet}{\Xi}$. In fact, we have $\Xi^{0}=\overset{\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$. Consequently,
@@
We set $E_{ij}^{\theta}=E_{ij}+E_{N+1-i,N+1-j},$ where $E_{ij}$ is the $N\times N$ matrix whose $(i,j)$-entry is $1$ and all other entries are 0. Let $e_{A}=f_{{\mathcal{O}}_{A}}$ be the characteristic function of the ${\mathrm{O}_{2r}}(q)$-orbit corresponding to $A\in\Xi_{N,2r}$. Therefore, $\{e_{A}|A\in\Xi_{N,2r}\}$ forms a basis of $S^{\jmath}(n,r)$. For convenience, set $e_{A}=0,\text{ if }A\notin\Xi_{N,2r}.$
@@
    $e_{B}*e_{A}=\sum_{p\in[1,N],a_{h+1,p}>0}{\boldsymbol{q}}^{\sum_{j>p}a_{hj}}\left[\!\left[a_{h,p}+1\right]\!\right]e_{{}_{h}\!A_{p}}.$
@@
$g_{h,A,p}=\begin{cases}{\boldsymbol{q}}^{\sum_{j<p}a_{h+1,j}}\left[\!\left[a_{h+1,p}+1\right]\!\right],&\text{ if }a_{h+1,p}>0;\\ 0,&\text{ if }a_{h+1,p}=0.\end{cases}$ (6.0.2)
@@
Observe from (4.6.1) that, if an ${\text{\rm O}}_{2r}$-orbit ${\mathcal{O}}_{A}$ in ${\mathcal{X}}\times{\mathcal{X}}$ splits into two ${\text{\rm SO}}_{2r}$-orbits $\check{\mathcal{O}}({}^{+}\!A^{\epsilon})\sqcup\check{\mathcal{O}}({}^{-}\!A^{-{\epsilon}})$, then the orbit function $f_{A}=f_{{}^{+}\!A^{\epsilon}}+f_{{}^{-}\!A^{-{\epsilon}}}\in\mathfrak{F}_{\check{G}}({\mathcal{X}}\times{\mathcal{X}})$. Thus, we have a $\mathbb{Z}$-algebra embedding $\mathfrak{F}_{G}({\mathcal{X}}\times{\mathcal{X}})\subset\mathfrak{F}_{\check{G}}({\mathcal{X}}\times{\mathcal{X}})$ which is the specialisation of the ${\mathcal{A}}$-algebra embedding $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)\subset S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ by Theorem 3.10. Composing with the isomorphism in Proposition 4.2 gives an algebra embedding $\iota:S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)\hookrightarrow S_{{\boldsymbol{q}}}^{\textsc{d}}(n,r)$. We make the convention that, when we write $e_{A}=\phi_{{}^{+}\!A^{\epsilon}}+\phi_{{}^{-}\!A^{-{\epsilon}}}$ for ${\epsilon}\in\{+,-\}$, we mean that $\iota(e_{A})=\phi_{{}^{+}\!A^{\epsilon}}+\phi_{{}^{-}\!A^{-{\epsilon}}}$. Here and below we use the sign convention
@@
$(1)\qquad{\mathcal{O}}_{A}$ $=\check{\mathcal{O}}({}^{+}\!A^{\epsilon})\sqcup\check{\mathcal{O}}({}^{-}\!A^{-{\epsilon}})$ (6.1.3)
@@ 6.1. The $B\in\overset{\bullet}{\Xi}$ case
By the claim, we see that $\phi_{\dot{B}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ (resp., $\phi_{\dot{B}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$) is a linear combination of $\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}$ (resp., $\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}$) if $h<n$, or a linear combination of $\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}$ and $\phi_{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}$ (resp., $\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}$ and $\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}$) if $h=n$. Note that, as standard basis elements in $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ defined by distinct $({\check{W}}_{\nu^{\bullet}},{\check{W}}_{\nu^{\bullet}})$ double cosets (see Definition 4.11(4)), $\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}$ and $\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}$ (or $\{\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}},\phi_{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}\}$ and $\{\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}},\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}\}$) are linearly independent. Hence, equating with (6.2.2) gives (6.2.1).
@@ 6.2. The case $B\in\overset{\circ\!\circ\!\circ}{\Xi}$ ($\implies h<n$)
Clearly, LHS$=\phi_{{}^{+}{B}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{B}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ since $\phi_{{}^{+}{B}^{+}}*\phi_{{}^{-}{A}^{-{\epsilon}}}=0=\phi_{{}^{-}{B}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}$. Also, every standard basis element $\phi_{{\check{\mathbb{A}}}^{\prime}}$ on the right hand side has row weight $\text{\rm rw}({\check{\mathbb{A}}}^{\prime})=\lambda^{+}$ or $\lambda^{-}$. Thus, (6.3.1) is obtained by multiplying idempotents $1_{\lambda^{+}}$ or $1_{\lambda^{-}}$ to the left hand side of (6.3.2). ∎
@@ 6.3. The case $B\in\overset{\circ\!\circ\!\bullet}{\Xi}$($\implies h=n$)
(1) If $A\in\overset{\bullet}{\Xi}$, then $\text{ro}(A)_{n+1}\neq 0$ and $(\text{\rm rw}(\dot{A}),\text{\rm cw}(\dot{A}))=(\mu^{\bullet},\nu^{\bullet}).$ Since $\lambda_{n+1}=0$ and $\text{ro}(A)_{n+1}=\text{co}(B)_{n+1}=2$, it follows that $a_{n+1,n+1}=2$ and $a_{n+1,p}=0$, for all $p\neq n+1$. Thus, the $(n+1,p)$ entry of ${}_{h}\!A_{p}$ if negative, for all $p\neq n+1$. Hence, (6.0.1) becomes $e_{{B}}*e_{A}=g_{n,A,{n+1}}e_{{{}_{n}A_{n+1}}}$. Since the central entry of ${}_{n}A_{n+1}$ is 0, we have ${\mathcal{O}}_{{}_{n}A_{n+1}}=\check{\mathcal{O}}(^{+}{{}_{n}A_{n+1}}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{{}_{n}A_{n+1}}^{-{\epsilon}})$ splits into two orbits and $\text{sgn}({{}_{n}A_{n+1}}^{\text{\normalsize$\llcorner$}})=\text{sgn}(A^{\llcorner})$, we have in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$,
@@
$(1)\quad\mathcal{O}_{A}$ $=\check{\mathcal{O}}(^{+}{A}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{A}^{-{\epsilon}})$ (7.0.4)
@@
We now follow the notaiton used in Proposition 3.8. Thus, $F_{i}=\langle v_{1},\ldots,v_{\widetilde{\alpha}_{i}}\rangle$ $(\alpha={\widehat{\lambda}})$ and $F^{\prime}=F^{A^{\prime}}$, where $A^{\prime}=_{n}\!A_{\overline{n+1}}$ has central entry 2. For $E\in{\mathcal{F}}$, by Corollary 5.3, we have $E_{n}=F_{n}+\langle v\rangle$, where $v$ has the form $v=c_{1}v_{\widetilde{\lambda}_{n}+1}+\cdots+c_{m}v_{\widetilde{\lambda}_{n}+m}+c_{r}v_{r}+c_{r+1}v_{r+1}$ with $m=a_{n+1,1}+\cdots+a_{n+1,n}$ and either $c_{r}=0$ or $c_{r+1}=0$, but not both 0.
@@
Clearly, LHS=$\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ since $\phi_{{}^{+}{C}^{+}}*\phi_{{}^{-}{A}^{{\epsilon}}}=\phi_{{}^{-}{C}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}=0.$ Moreover, every standard basis element $\phi_{\check{\mathbb{A}}^{\prime}}$ on the right hand side satisfies cw$(\check{\mathbb{A}}^{\prime})=\nu^{\epsilon}$ or $\nu^{-{\epsilon}}$. Multiplying idempotents $1_{\nu^{\epsilon}}(\text{resp., }1_{\nu^{-{\epsilon}}})$ to the right hand sides of (7.3.1) yields the required formulas for $\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}(\text{resp., }\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}})$.

### v2

@@ 4.3. The geometric definition of the natural basis
(2) If $A\in\overset{\circ\!\circ\!\bullet}{\Xi}$ then, by Corollary 2.7, $W_{\lambda}={\check{W}}_{\lambda^{+}}$ and ${\check{W}}_{\mu}={\check{W}}_{\mu^{\bullet}}$,
@@
If $d_{A}s_{r}=s_{r}d_{A}$ then, by Lemma 2.2 (2), $a_{n+1,n+1}>0$. Thus, $A\in\overset{\bullet}{\Xi}$ and one of the two disjoint sets in (4.12.1) is a double coset in ${\check{W}}$, depending on $d_{A}\in{\check{W}}$ or not.
@@
Recalled the natural basis $\{\phi_{\alpha,\beta}^{d}\mid(\alpha,d,\beta)\in\check{\mathcal{D}}(n,r)\}$ for $S_{{\boldsymbol{q}}}^{\textsc{d}}(n,r)$ given in Corollary 4.3. With the bijection $\eta$ in Lemma 4.12, we may also label the basis element by signed matrix notation:
@@
(1) For $A\in\overset{\circ\!\circ\!\circ}{\Xi}$ and ${\epsilon}\in\{+,-\}$, if $A$ defines ${}^{+}{A}^{{\epsilon}},^{-}{A}^{-{\epsilon}}\in\overset{\circ\!\circ\!\circ}{\Xi}_{+}\sqcup\overset{\circ\!\circ\!\circ}{\Xi}_{-}$, then $\eta(^{+}{A}^{{\epsilon}})=(\lambda^{+},d_{1},\mu^{\epsilon})$ and $\eta(^{-}{A}^{-{\epsilon}})=(\lambda^{-},d_{2},\mu^{-{\epsilon}})$ (see (6.1.1) for notational convention and Definition 4.11 for the definition of $d_{1},d_{2}$) which define distinct natural basis elements $\phi_{{}^{+}{A}^{{\epsilon}}}$ and $\phi_{{}^{-}{A}^{-{\epsilon}}}$ according their row or column weights given in (4.12.5). This distinction can be reflected by the distinction of the ${\mathsf{m}}_{1}\times{\mathsf{m}}_{2}$ subset matrices $\big{(}{\check{R}}_{i}^{\lambda^{+}}\cap d_{1}{\check{R}}_{j}^{\mu^{\epsilon}}\big{)}$ and $\big{(}{\check{R}}_{i}^{\lambda^{-}}\cap d_{2}{\check{R}}_{j}^{\mu^{-{\epsilon}}}\big{)}$ since
@@
(2) For $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$, we also have ${}^{+}{A}^{{\epsilon}},^{-}{A}^{-{\epsilon}}\in\overset{\bullet\!\circ\!\bullet}{\Xi}_{+}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}_{-}$. In this case, the natural basis elements $\phi_{{}^{+}{A}^{{\epsilon}}}$ and $\phi_{{}^{-}{A}^{-{\epsilon}}}$ have the same domain and codomain, but are defined by distinct double cosets ${\check{W}}_{\lambda^{\bullet}}d_{1}{\check{W}}_{\mu^{\bullet}}$ and ${\check{W}}_{\lambda^{\bullet}}d_{2}{\check{W}}_{\mu^{\bullet}}$. Of course, the conjugate intersections ${\check{W}}_{\lambda^{\bullet}}\cap({\check{W}}_{\mu^{\bullet}})^{d_{1}^{-1}}$ and ${\check{W}}_{\lambda^{\bullet}}\cap({\check{W}}_{\mu^{\bullet}})^{d_{2}^{-1}}$ are also distinct in this case.
@@
(1). The theorem shows that the ${\mathcal{A}}$-algebra $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ is the “quantumization” (in the sense of [8, p.17]) of the algebras $\{\mathfrak{F}_{{\mathrm{SO}}_{2r}(q)}({\mathcal{X}}\times{\mathcal{X}})\}_{q\in\mathcal{P}}$ which is the same as the algebra defined in [14, (5)]. It should be pointed out that the dimension formula #$\#\Xi_{\textsc{d}}$ given in [14, Lem. 4.2.1] is incorrect as the definition of the index set $\Xi_{\textsc{d}}=\Xi^{+}\sqcup\Xi^{0}\sqcup\Xi^{-}$ mistakenly used $\Xi^{0}$ instead of $\overset{\bullet}{\Xi}$. In fact, we have $\Xi^{0}=\overset{\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$. Consequently,
@@
(2). The anonymous referee pointed out to us that by using coordinate algebra approach developed in [18], the dimension formula (4.15.1) of $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ was also obtained by Ziqing Xiang in an unpublished manuscript.
@@
Let $e_{A}=f_{{\mathcal{O}}_{A}}$ be the characteristic function of the ${\mathrm{O}_{2r}}(q)$-orbit corresponding to $A\in\Xi_{N,2r}$. Therefore, $\{e_{A}|A\in\Xi_{N,2r}\}$ forms a basis of $S^{\jmath}(n,r)$. For convenience, set $e_{A}=0,\text{ if }A\notin\Xi_{N,2r}.$
@@
    $e_{B}*e_{A}=\sum_{p\in[1,N],a_{h+1,p}>0}{\boldsymbol{q}}^{\sum_{j>p}a_{h,j}}\left[\!\left[a_{h,p}+1\right]\!\right]e_{{}_{h}\!A_{p}}.$
@@
$g_{h,A,p}=\begin{cases}{\boldsymbol{q}}^{\sum_{j>p}a_{h,j}}\left[\!\left[a_{h,p}+1\right]\!\right],&\text{ if }a_{h+1,p}>0;\\ 0,&\text{ if }a_{h+1,p}=0.\end{cases}$ (6.0.2)
@@
Observe from (4.6.1) that, if an ${\text{\rm O}}_{2r}$-orbit ${\mathcal{O}}_{A}$ in ${\mathcal{X}}\times{\mathcal{X}}$ splits into two ${\text{\rm SO}}_{2r}$-orbits $\check{\mathcal{O}}({}^{+}\!A^{\epsilon})\sqcup\check{\mathcal{O}}({}^{-}\!A^{-{\epsilon}})$, then the orbit function $f_{A}=f_{{}^{+}\!A^{\epsilon}}+f_{{}^{-}\!A^{-{\epsilon}}}\in\mathfrak{F}_{\check{G}}({\mathcal{X}}\times{\mathcal{X}})$. Thus, we have a $\mathbb{Z}$-algebra embedding $\mathfrak{F}_{G}({\mathcal{X}}\times{\mathcal{X}})\subset\mathfrak{F}_{\check{G}}({\mathcal{X}}\times{\mathcal{X}})$ which is the specialization of the ${\mathcal{A}}$-algebra embedding $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)\subset S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ by Theorem 3.10. Composing with the isomorphism in Proposition 4.2 gives an algebra embedding $\iota:S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)\hookrightarrow S_{{\boldsymbol{q}}}^{\textsc{d}}(n,r)$. We make the convention that, when we write $e_{A}=\phi_{{}^{+}\!A^{\epsilon}}+\phi_{{}^{-}\!A^{-{\epsilon}}}$ for ${\epsilon}\in\{+,-\}$, we mean that $\iota(e_{A})=\phi_{{}^{+}\!A^{\epsilon}}+\phi_{{}^{-}\!A^{-{\epsilon}}}$. Here and below we use the sign convention
@@
$(1)\qquad{\mathcal{O}}_{A}$ $=\check{\mathcal{O}}({}^{+}\!A^{\epsilon})\sqcup\check{\mathcal{O}}({}^{-}\!A^{-{\epsilon}}),$ (6.1.3)
@@ 6.1. The $B\in\overset{\bullet}{\Xi}$ case
By the claim, we see that $\phi_{\dot{B}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ (resp., $\phi_{\dot{B}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$) is a linear combination of $\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}$ (resp., $\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}$) if $h<n$, or a linear combination of $\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}$ and $\phi_{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}$ (resp., $\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}$ and $\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}$) if $h=n$. Note that, as natural basis elements in $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ defined by distinct $({\check{W}}_{\lambda^{\bullet}},{\check{W}}_{\mu^{\bullet}})$ double cosets (see Definition 4.11(4)), $\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}$ and $\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}$ (or $\{\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}},\phi_{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}\}$ and $\{\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}},\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}\}$) are linearly independent. Hence, equating with (6.2.2) gives (6.2.1).
@@ 6.2. The case $B\in\overset{\circ\!\circ\!\circ}{\Xi}$ ($\implies h<n$)
Clearly, LHS$=\phi_{{}^{+}{B}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{B}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ since $\phi_{{}^{+}{B}^{+}}*\phi_{{}^{-}{A}^{-{\epsilon}}}=0=\phi_{{}^{-}{B}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}$. Also, every natural basis element $\phi_{{\check{\mathbb{A}}}^{\prime}}$ on the right hand side has row weight $\text{\rm rw}({\check{\mathbb{A}}}^{\prime})=\lambda^{+}$ or $\lambda^{-}$. Thus, (6.3.1) is obtained by multiplying idempotents $1_{\lambda^{+}}$ or $1_{\lambda^{-}}$ to the left hand side of (6.3.2). ∎
@@ 6.3. The case $B\in\overset{\circ\!\circ\!\bullet}{\Xi}$($\implies h=n$)
(1) If $A\in\overset{\bullet}{\Xi}$, then $\text{ro}(A)_{n+1}\neq 0$ and $(\text{\rm rw}(\dot{A}),\text{\rm cw}(\dot{A}))=(\mu^{\bullet},\nu^{\bullet}).$ Since $\lambda_{n+1}=0$ and $\text{ro}(A)_{n+1}=\text{co}(B)_{n+1}=2$, it follows that $a_{n+1,n+1}=2$ and $a_{n+1,p}=0$, for all $p\neq n+1$. Thus, the $(n+1,p)$ entry of ${}_{h}\!A_{p}$ is negative, for all $p\neq n+1$. Hence, (6.0.1) becomes $e_{{B}}*e_{A}=g_{n,A,{n+1}}e_{{{}_{n}A_{n+1}}}$. Since the central entry of ${}_{n}A_{n+1}$ is 0, we have ${\mathcal{O}}_{{}_{n}A_{n+1}}=\check{\mathcal{O}}(^{+}{{}_{n}A_{n+1}}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{{}_{n}A_{n+1}}^{-{\epsilon}})$ splits into two orbits and $\text{sgn}({{}_{n}A_{n+1}}^{\text{\normalsize$\llcorner$}})=\text{sgn}(A^{\llcorner})$, we have in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$,
@@
$(1)\quad\mathcal{O}_{A}$ $=\check{\mathcal{O}}(^{+}{A}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{A}^{-{\epsilon}}),$ (7.0.4)
@@
We now follow the notation used in Proposition 3.8. Thus, $F_{i}=\langle v_{1},\ldots,v_{\widetilde{\alpha}_{i}}\rangle$ $(\alpha={\widehat{\lambda}})$ and $F^{\prime}=F^{A^{\prime}}$, where $A^{\prime}=_{n}\!A_{\overline{n+1}}$ has central entry 2. For $E\in{\mathcal{F}}$, by Corollary 5.3, we have $E_{n}=F_{n}+\langle v\rangle$, where $v$ has the form $v=c_{1}v_{\widetilde{\lambda}_{n}+1}+\cdots+c_{m}v_{\widetilde{\lambda}_{n}+m}+c_{r}v_{r}+c_{r+1}v_{r+1}$ with $m=a_{n+1,1}+\cdots+a_{n+1,n}$ and either $c_{r}=0$ or $c_{r+1}=0$, but not both 0.
@@
Clearly, LHS=$\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ since $\phi_{{}^{+}{C}^{+}}*\phi_{{}^{-}{A}^{{\epsilon}}}=\phi_{{}^{-}{C}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}=0.$ Moreover, every natural basis element $\phi_{\check{\mathbb{A}}^{\prime}}$ on the right hand side satisfies cw$(\check{\mathbb{A}}^{\prime})=\nu^{\epsilon}$ or $\nu^{-{\epsilon}}$. Multiplying idempotents $1_{\nu^{\epsilon}}(\text{resp., }1_{\nu^{-{\epsilon}}})$ to the right hand sides of (7.3.1) yields the required formulas for $\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}(\text{resp., }\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}})$.